### v5

@@ 6 Numerical results
All of our algorithms were written [^1] The code is available on the arXiv page of this paper and on https://github.com/melaniesi/EdgeExpansion.jl in Julia [6] version 1.9.2. That is, the split-and-bound Algorithm 2 including pre-elimination and the transformation from $k$-bisection to max-cut problems. Also, Algorithm 3 we implemented in Julia. The SDPs to compute our cheap lower bounds $\ell_{k}$ from the bisection problem in 7 are solved with MOSEK 10.0 [30] using JuMP [26]. We also use JuMP to solve MIQCPs with Gurobi [15] version 11.0. The solver BiqBin [16] for binary quadratic problems was used to solve the parametrized problems in Dinkelbach’s method, and we extended the C code of this solver by adding the option to provide an initial lower bound on the maximization problem. The corresponding changes are tracked in the git repository https://gitlab.aau.at/BiqBin/biqbin. All computations were carried out on an AMD EPYC 7532 with 32 cores with 3.30GHz and 1024GB RAM, operated under Debian GNU/Linux 11.
@@ 6.2 Discussion of the experiments
In the section for Dinkelbach’s algorithm, the first column gives the first guess for the edge expansion, i.e., the first trial for $\gamma$. As described before, we take the upper bound from the heuristic for this initialization. Note, that this first guess may differ from $u^{*}$, since in the pre-elimination step of split & bound we perform 30 additional rounds of simulated annealing for all indices $k\in\mathcal{I}$. Column 2 indicates how many parametrized problems $P_{\gamma_{i}}$ have been solved, and column 3 gives the total number of branch-and-bound nodes for solving all parametrized problems. The fourth column of the results of Dinkelbach’s algorithm displays the total time, including running the heuristic to obtain the first guess.
@@
As can be seen in all tables, the pre-elimination phase of split & bound only leaves a comparably small number of candidates for $k$ to be further investigated. Remember that the number of potential candidates is $\lfloor\frac{n}{2}\rfloor$, whereas $|\mathcal{I}|$ is the number of candidates that remain after the pre-elimination. For 12 instances we were able to compute the edge expansion even within the pre-elimination phase, i.e. the set of candidates $\mathcal{I}$ was empty. For the randomly generated 0/1-polytopes on average only 12% of the candidates have to be further examined, for the other instance classes we can approximately eliminate 80% of the candidate values for $k$ on average. Only for 6 instances we were not able to half the number of candidates within the pre-elimination phase. All of those 6 instances are from the DIMACS or network graphs test set. This indicates that in general already the cheap SDP bound is of good quality.
@@
Whenever the heuristic already returns the value of the optimal solution, we only have to solve one parametrized problem to certify the optimality of this value. For most of the instances tested, this certificate is already obtained in the root node of the branch-and-bound tree. However, there are many instances where BiqBin terminates because of numerical problems even for the first parametrized problem, see § 6.2. This in particular arises when $\gamma_{d}$ and $\gamma_{n}$ (see Section § 5.1) are large.
@@ Conclusion
Table 2: Comparison of Algorithm 2 (split & bound), Algorithm 3 (Dinkelbach) and Gurobi for graphs of random 0/1-polytopes. The last column displays the gap reported by Gurobi after a time limit of 3 hours. * Increased timelimit of Gurobi to match the time one of the other algorithms needed for exact computation.
@@
00100100100100200200200200300300300300400400400400500500500500000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111time (sec)of instances solvedGurobiDinkelbachSplit & Bound
@@ Appendix A Computing the maximum cut of a graph
$\max_{X}\,\Big{\{}\frac{1}{4}\langle L,X\rangle:\operatorname{diag}(X)=e,\,\mathcal{A}(X)=b,\,X\succeq 0\Big{\}}$ (12)
@@
$f(\gamma)$ $=\max_{X}\Big{\{}\frac{1}{4}\langle L,X\rangle-\gamma^{\top}(\mathcal{A}(X)-b):\operatorname{diag}(X)\ =\ e,\,X\succeq 0\Big{\}}$
$=b^{\top}\gamma+\max_{X}\Big{\{}\Big{\langle}\frac{1}{4}L-\mathcal{A}^{\top}(\gamma),X\Big{\rangle}:\operatorname{diag}(X)=e,\,X\succeq 0\Big{\}}$
@@
$\min_{\gamma}\{f(\gamma):\gamma\geq 0\},$

### v6

@@ 6 Numerical results
All of our algorithms were written [^1] The code is available on the arXiv page https://arxiv.org/abs/2403.04657 and on https://github.com/melaniesi/EdgeExpansion.jl in Julia [6] version 1.9.2. That is, the split-and-bound Algorithm 2 including pre-elimination and the transformation from $k$-bisection to max-cut problems. Also, Algorithm 3 we implemented in Julia. The SDPs to compute our cheap lower bounds $\ell_{k}$ from the bisection problem in 7 are solved with MOSEK 10.0 [31] using JuMP [27]. We also use JuMP to solve MIQCPs with Gurobi [16] version 11.0. The solver BiqBin [17] for binary quadratic problems was used to solve the parametrized problems in Dinkelbach’s method, and we extended the C code of this solver by adding the option to provide an initial lower bound on the maximization problem. The corresponding changes are tracked in the git repository https://gitlab.aau.at/BiqBin/biqbin. All computations were carried out on an AMD EPYC 7532 with 32 cores with 3.30GHz and 1024GB RAM, operated under Debian GNU/Linux 11.
@@ 6.2 Discussion of the experiments
In the section for Dinkelbach’s algorithm, the first column gives the first guess for the edge expansion, i.e., the first trial for $\gamma$. As described before, we take the upper bound from the heuristic for this initialization. Note, that this first guess may differ from $u^{*}$, since in the pre-elimination step of split & bound we perform 30 additional rounds of simulated annealing for all indices $k\in\mathcal{I}$. Column 2 indicates how many parametrized problems $P(\gamma_{i})$ have been solved, and column 3 gives the total number of branch-and-bound nodes for solving all parametrized problems. The fourth column of the results of Dinkelbach’s algorithm displays the total time, including running the heuristic to obtain the first guess.
@@
As can be seen in all tables, the pre-elimination phase of split & bound only leaves a comparably small number of candidates for $k$ to be further investigated. Remember that the number of potential candidates is $\lfloor\frac{n}{2}\rfloor$, whereas $|\mathcal{I}|$ is the number of candidates that remain after the pre-elimination. For 12 instances we were able to compute the edge expansion even within the pre-elimination phase, i.e. the set of candidates $\mathcal{I}$ was empty. For the randomly generated 0/1-polytopes on average only 12% of the candidates have to be further examined, for the other instance classes we can approximately eliminate 80% of the candidate values for $k$ on average. Only for 6 instances we were not able to halve the number of candidates within the pre-elimination phase. All of those 6 instances are from the DIMACS or network graphs test set. This indicates that in general already the cheap SDP bound is of good quality.
@@
Whenever the heuristic already returns the value of the optimal solution, we only have to solve one parametrized problem to certify the optimality of this value. For most of the instances tested, this certificate is already obtained in the root node of the branch-and-bound tree. However, there are many instances among the random 0/1 polytopes where BiqBin terminates because of numerical problems even for the first parametrized problem, see § 6.2. This in particular arises when $\gamma_{d}$ and $\gamma_{n}$ (see Section § 5.1) are large.
@@ Conclusion
Table 2: Comparison of Algorithm 2 (split & bound), Algorithm 3 (Dinkelbach) and Gurobi for graphs of random 0/1-polytopes. A “-” indicates that Algorithm 3 could not solve this instance. The last column displays the gap reported by Gurobi after a time limit of 3 hours. * Increased timelimit of Gurobi to match the time one of the other algorithms needed for exact computation.
@@
010010020020030030040040050050000.20.20.40.40.60.60.80.811time (sec)% of instances solvedGurobiDinkelbachSplit & Bound
@@ Appendix A Computing the maximum cut of a graph
$\max_{X}\,\Big\{\frac{1}{4}\langle L,X\rangle\mathrel{\mathop{\ordinarycolon}}\operatorname{diag}(X)=e,\,\mathcal{A}(X)=b,\,X\succeq 0\Big\}$ (12)
@@
$f(\gamma)$ $=\max_{X}\Big\{\frac{1}{4}\langle L,X\rangle-\gamma^{\top}(\mathcal{A}(X)-b)\mathrel{\mathop{\ordinarycolon}}\operatorname{diag}(X)\ =\ e,\,X\succeq 0\Big\}$
$=b^{\top}\gamma+\max_{X}\Big\{\Big\langle\frac{1}{4}L-\mathcal{A}^{\top}(\gamma),X\Big\rangle\mathrel{\mathop{\ordinarycolon}}\operatorname{diag}(X)=e,\,X\succeq 0\Big\}$
@@
$\min_{\gamma}\{f(\gamma)\mathrel{\mathop{\ordinarycolon}}\gamma\geq 0\},$